Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 572
Entry time: 09 Dec 2024, 09:56
|
Author: |
Frederik Wolff-Fabris |
Location: |
General |
Severity: |
Information |
Subject: |
Notes XO Integration Meeting 09.12.2024 |
|
|
General Announcement:
- A. Galler informs two downtimes happened last two nights totalizing 10-12hs due to communication issues (between terminals) at the accelerator.
CTRL:
- Nothing to report.
DA:
- Not present.
EEE:
- Nothing to report in issues.
- Tomorrow from 12:30-13:30 a readout of parameters in tunnels will be made; PLC updates has also a plan.
VAC:
- Everything fine with vacuum.
XPD:
- Reminder the photon ramp-down days take place tomorrow; SA3 hutches must be interlocked.
- SA2 XTD1 XGM stopped working and BKR could resolve by experienced crew.
DET:
- Nothing major to report.
ITDM:
- Nothing to report.
XRO:
- M1 in SA2: a ZZ on Friday take place to switch OFF and disconnect a controller that was misbehaving and was acting on the piezo without Karabo control.
- U. Boesenberg reports it got better on Saturday and it did not completely fixed the source of problem (beam jitter). Discussion with Johannes M. and Ulrike B. will take place.
LAS:
- Nothing to report.
UNSYS:
- PS cell20 SA3 to be fixed in the WMP.
SPB:
- Nothing to report.
- Request for tuning to improve photon intensities higher than 1mJ today. To be addressed by XO and PRC.
FXE:
- Nothing to report.
MID:
- Apart of beam stability on Friday; a radiation Interlock was triggered by a Lamp. As consequence, could not interlock the hutch. Got solved temporarily and further investigation on the WMP.
HED:
- Not present.
SCS:
- Not present.
SQS:
- Nothing to report.
SXP:
- Not present. |